OpenStudy (anonymous):

what is the slope of a line perpendicular to x-y=16

OpenStudy (jdoe0001):

what's the slope of x-y=16?

OpenStudy (anonymous):

idk

OpenStudy (jdoe0001):

well, solve for "y", what do you get?

jimthompson5910 (jim_thompson5910):

solve for y to get x-y=16 -y=16 - x -1*(-y)=-1*(16 - x) y = -16 + x y = x - 16 what is the slope of this line?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

correct, the slope is 1

jimthompson5910 (jim_thompson5910):

1 = 1/1 so what you do is you flip the fraction and you flip the sign to get -1/1 = -1

jimthompson5910 (jim_thompson5910):

therefore, the perpendicular slope is -1

OpenStudy (anonymous):

ok what is the slope of a line perpendicular to x-y=16

jimthompson5910 (jim_thompson5910):

-1, I just wrote it above

jimthompson5910 (jim_thompson5910):

and all the steps are above as well

OpenStudy (anonymous):

wrong ? sorry hold on

jimthompson5910 (jim_thompson5910):

well it's -1, there must be a typo somewhere

OpenStudy (anonymous):

what is y=-2/3x+4 in standard form

jimthompson5910 (jim_thompson5910):

I'll start you off y=-2/3x+4 3y=-2x+12 .. multiply EVERY term by the LCD 3 to clear out the fraction I'll let you take over

OpenStudy (anonymous):

well 3y/3 is y and -2/3 is -0.66 and 12/3 is 4

OpenStudy (anonymous):

is that right

jimthompson5910 (jim_thompson5910):

here, you add 2x to both sides to get this 3y=-2x+12 3y+2x = 12 2x + 3y = 12

jimthompson5910 (jim_thompson5910):

so y=-2/3x+4 in standard form is 2x + 3y = 12

OpenStudy (anonymous):

o thanks so much

jimthompson5910 (jim_thompson5910):

yw